The short run production function Q=L^3-2L^2-4L+40. determine the amount of laborers



1
Expert's answer
2021-07-30T11:47:01-0400

"Q=L^3-2L^2-4L+40"

"\frac{\delta Q}{\delta L} =3L^2-4L-4"

Equate "\frac{\delta Q}{\delta L}=0"

"\implies" "3L^2-4L-4=0"

Solving the above quadratic equation:


"L= \frac{4+(4^2-[4\times3\times-4])^\frac {1}{2}}{2\times3}"

"L=2" labourers